Monthy Proportion of On Tme Payments
80
82
84
83
83
84
85
84
82
83
84
83

Calculate the following for a 3-month moving average forecast; round each to 2 decimal places:

a. MFE

b. MAE

c. MSE

d. MAPE
Answer #1

The following is the data provided: Period Data January 80 February March April May June July August September October Novemb

The 3-month moving average forecasts are calculated in the table below: Period Data MA(3) Forecasts January February March Ap

Now, the following table shows the calculations of the corresponding error metrics: Period MA(3) Forecasts Abs. Error Error A

October 83.6667 83 - 83.6667 = 0.6667 (83 - 83.6667)2 = 0.4444 83-83.6667 83 0.8% 84 - 83 = 84-83 November 84 (84 - 83)2 = 1

Monthy Proportion of On Tme Payments (A) Forecast (F) A-F Abs (A-F)
80
82
84
83 82 1 1
83 83 0 0
84 83.3333 0.6667 0.6667
85 83.3333 1.6667 1.6667
84 84 0 0
82 84.3333 -2.3333 2.3333
83 83.6667 -0.6667 0.6667
84 83 1 1
83 83 0 0
Sum 1.3334 7.3334

MFE = sum of (A-F) / n = 1.3334/12 = 0.11

MAE =Mean Abs of (A-F) = 7.3334 / 9 = 0.82